An air conditioner (AC) removes heat at a rate of 1.2kJ/sec from a room. A power of 400W is required to run the AC . The coefficient of performance of the AC is 10% of that of the refrigerator operating between outside and room temperature. If outside temperature is 37°C what will be room temperature?
Ans : 27°C
1
Expert's answer
2020-11-23T10:25:36-0500

As per the given question,

Rate of removal of heat =1.2KJ/s = 1200J/s

Power required to run AC = 400W

Coefficient of performance of AC=10%

Outside temperature "T_H=37^\circ"

Room temperature "T_c=?"

We know that coefficient of performance"=1-\frac{T_c}{T_H}"

Now, substituting the values,

"0.1=1-\frac{T_c}{37}"

"\Rightarrow \frac{T_c}{37}=0.9"

"\Rightarrow T_c=0.9\times 37=33.3^\circ"